A 0.460 inch diameter brass rod is 8'5 7/32 long. Round answers to A. and B. to 3 significant digits.
A. Find the volume in cubic inches, of the rod.
B.Compute the total weight of 40 rods. Brass weighs 0.300 pound per cubic inch.

Answer by nerdybill(7384)

.
A. Find the volume in cubic inches, of the rod.
Volume = ("area of base")* height
.
Since the diameter is given as: 0.460 inch
The radius is 1/2(diameter) = 0.230 inch
.
"area of base" = (pi)r^2 = (3.142)(0.230)^2 = (3.142)(0.0529) = 0.166 sq inches
"height" = 8'5 7/32 = (8*12)+5+(7/32) = 96+5+(0.21875) = 101.21875 inches
Volume then is:
0.166 * 101.21875 = 16.802 cubic inches
.
B.Compute the total weight of 40 rods. Brass weighs 0.300 pound per cubic inch
That would be:
40 * 16.802 * 0.300 = 201.624 pounds